OpenStudy (anonymous):

Perimeter of one triangle is 7/10 the perimeter of the a similar triangle. The difference of two corresponding sides is 3in. What are the lengths of the sides?

OpenStudy (anonymous):

@amistre64 @RadEn @Hero @cwrw238 @tcarroll010 please help

OpenStudy (anonymous):

If x = the side of one triangle and y = the corresponding side of the other and similar triangle, then: x = (7/10)y Also, x = y - 3 You now have 2 equations in 2 variables that can easily be solved by substitution. Show your work and I'll help you further if you get stuck.

OpenStudy (anonymous):

y-3 = 7/10y -3/10y = -3 y = 10 and x = 7

OpenStudy (anonymous):

There you go! You got it!

OpenStudy (anonymous):

The reason you can extend this logic from perimeter to side length is that each ot the corresponding sides are 7/10 of the other. The perimeter was just a device to talk about the sides and one pair of sides in particular.

OpenStudy (anonymous):

@tcarroll010 Thanks for the explanation!

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @vmathhelp

OpenStudy (anonymous):

You're very welcome!